OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 29°, b = 26, c = 28 A = 92.5°, C = 58.5°, a = 53.6; A = 87.5°, C = 121.5°, a = 53.6 A = 92.5°, C = 58.5°, a = 12.6; A = 87.5°, C = 121.5°, a = 12.6 A = 119.5°, C = 31.5°, a = 46.7; A = 2.5°, C = 148.5°, a = 2.3 A = 119.5°, C = 31.5°, a = 14.5; A = 2.5°, C = 148.5°, a = 14.5

OpenStudy (anonymous):

1st two are out it's 3 or 4

OpenStudy (anonymous):

\[\frac{ \sin A }{ a }=\frac{ \sin B }{b }=\frac{ \sin C }{ c}\] You have enough information there to fill in 3 out of 4 of the parts.